OpenStudy (anonymous):

show that if you apply a particular factoring formula to 9x^2-4 that you get (3x-2)(3x+2) How do I do this?

OpenStudy (cwrw238):

this is the difference of two squares a^2 - b^2 = (a + b)(a - b)

OpenStudy (cwrw238):

so in your function a = 3x and b = 2

OpenStudy (anonymous):

work it backwards to show that (3x-2)(3x+2)=9x^2-4 FOIL (3x-2)(3x+2)

OpenStudy (anonymous):

I dont get it?

OpenStudy (anonymous):

I'll try it...

OpenStudy (anonymous):

always true? ?

OpenStudy (anonymous):

always true for difference of 2 squares

OpenStudy (anonymous):

Do i solve it as a equation?

OpenStudy (anonymous):

no, just FOIL (3x+2)(3x-2), you will end up with 9x^2-4 because the -6x and the +6x cancel

OpenStudy (cwrw238):

a^2 = 9x^2 so a = 3x b^1 = 4 so b = 2 now plug these values into the formula for a^2 - b^2

OpenStudy (anonymous):

I dont get it!!

OpenStudy (anonymous):

can someone explain this in a special ed term?

OpenStudy (anonymous):

\[3x^{2}-2^{2}=(3x+2)(3x-2)\]